OpenStudy (anonymous):

Problem check: How many 6 card hands containing no queens and no spades can be chosen from a bridge deck of 52 card? My attempt: I used the combination formula, but before that, I did 52-8 (4 because of the queens, another 4 for the spades). So I was left with 44!/(38!6!). My result from that...I dunno, I don't feel it is right or if what I was even so.

OpenStudy (anonymous):

@shamil98 Raito is a genius, no? :D

OpenStudy (anonymous):

Does the 6 not come in to play at all? May I ask how you figured?

OpenStudy (anonymous):

Oh sorry, I meant it should be \(\displaystyle\binom{52-8}{6}\). You're choosing only 6 of the total possible cards, which are the original 52, excluding the 4 aces and 4 queens.

OpenStudy (anonymous):

Right, and that notation indicates 44!/6!?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

God damn it I was right after I did it again. Thank you so much.

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

@SithsAndGiggles 3692043853872845511171702515298077625443942400000000 that does not make sense to me o.o did I do it wrong? Sorry once again o.o

OpenStudy (anonymous):

For anyone looking, help is appreciated.

OpenStudy (kropot72):

There are 4 queens (including the queen of spades) and 12 spade cards (excluding the queen of spades). Therefore the number of cards to choose from is 52 - 16 = 36. The possible number of combinations is \[36C6=\frac{36!}{6!30!}=\frac{36\times35\times34\times33\times32\times31}{6\times5\times4\times3\times2\times1}\]

OpenStudy (anonymous):

I can't believe I didn't see that...thank you so much.

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

I can't either. Thought it said aces, not spades... sorry!